OpenStudy (anonymous):

How to Solve? 1/4th root of (x-x^4)≥0 (Clarify: mean 1 over the 4th root of x-x^4)

OpenStudy (calculusfunctions):

\[\sqrt[4]{x -x ^{4}}\ge 0\]Correct?

OpenStudy (anonymous):

no, its that in the denominator with a 1 on top

OpenStudy (tkhunny):

First, think about Domain Issues. If \[x \le x^4\], there is no solution in the Real Numbers

OpenStudy (calculusfunctions):

\[\frac{ 1 }{ \sqrt[4]{x -x ^{4}} }\ge 0\]Correct?

OpenStudy (anonymous):

yes

OpenStudy (calculusfunctions):

Now first we must find the domain (restrictions on the variable). Do you know how to do that? What @tkhunny said is not true so don't be influenced by that.

OpenStudy (anonymous):

I'm not really sure how to

OpenStudy (calculusfunctions):

For example, the denominator cannot be zero and the radicand cannot be negative for even roots, if the solution is to be in the reals. Correct?

OpenStudy (calculusfunctions):

Radicand is the expression inside the root sign.

OpenStudy (anonymous):

Okay, i understand that

OpenStudy (calculusfunctions):

@Studentc14 do you understand?

OpenStudy (calculusfunctions):

Good!

OpenStudy (tkhunny):

Not sure how to respond to that. You call it incorrect when I say it but then repeat it yourself? Please rethink your disclaimer against my correct statement in the beginning.

OpenStudy (calculusfunctions):

So then can you tell me what the domain is?

OpenStudy (anonymous):

well i understand what it is but i dont know how to go about it with finding it... i know that 1 is a not a solution because that gives you 0

OpenStudy (calculusfunctions):

@tkhunny you said that there is no solution in the real numbers but if x = 0.1, for example, would the radicand not be a real number?

OpenStudy (anonymous):

if you wouldn't mind, just please ignore each other and help me solve the problem.

OpenStudy (tkhunny):

Incorrect. Go read it again. I was talking about the Domain, just like you. I was not offering a solution. Please read more carefully, certainly prior to correcting things that don't erquire correction. On the other hand, I do apprecieate your zeal in defending the student.

OpenStudy (calculusfunctions):

OK @Studentc14 I'll walk you through the steps, hence guiding you towars the correct solution, without giving away the answers. Alright? Ready?

OpenStudy (anonymous):

ok

OpenStudy (tkhunny):

@Studentc14 - learning to read and understand the writings of others is an important part of your learning. Anyway, I'll leave it to calculusfunctions from here. Good luck.

OpenStudy (calculusfunctions):

@tkhunny you said that\[x \le x ^{4}\]has no solution in the reals. Except that it does. For example, if x = 0.1 then\[x \le x ^{4}\]

OpenStudy (tkhunny):

Not so. You missed the "if" AND you failed to address the orignal problem. Please read the whole statement with reference to the original problem.

OpenStudy (calculusfunctions):

@Studentc14 first since the denominator cannot equal zero and the radical is in the denominator,\[x -x ^{4}>0\]Correct?

OpenStudy (anonymous):

yes

OpenStudy (calculusfunctions):

OK @Studentc14 then we first need to solve this inequality to find the domain. Do you know how to solve this inequality?

OpenStudy (anonymous):

not really..

OpenStudy (calculusfunctions):

OK first we factor the expression on the left side of the inequality. Can you do that?

OpenStudy (anonymous):

no i dont know how to do it with 4th powers

OpenStudy (calculusfunctions):

You're over thinking it. It doesn't matter what the powers are. Step on e is always to decide whether there is a greatest common factor. Is there a greatest common factor?

OpenStudy (anonymous):

x? then it would be x(1-x^3)

OpenStudy (calculusfunctions):

What is the greatest common factor of\[x -x ^{4}\]

OpenStudy (calculusfunctions):

Perfect!

OpenStudy (calculusfunctions):

So then we have\[x(1-x ^{3})>0\]Correct? FYI that's exponent 3. With the equation editor on this site, the exponents 3 and 2 look the same. As long as you know that's a 3. OK?

OpenStudy (calculusfunctions):

Since we're only interested to find x values in the reals  In other words in which of these interval(s) is\[x(1-x ^{3})>0\]Can you finish that interval line and tell me?